OpenStudy (anonymous):

Joe is 10 years older than James. In 8 years, twice Joe's age will equal three times James's age. How old is each now?

OpenStudy (anonymous):

x=joe's age y=james' age x=y+10 2(x+8)=3(y+10) 2x+16=3y+30 2(y+10)=3y+14 y=6 x=16